Indicate how the concentration of each species in the chemical equation will change to reestablish equilibrium after increasing the pressure. 

$$ \mathbf{N}_{2}(g)+3 \mathbf{H}_{2}(g) \rightleftharpoons 2 \mathrm{NH}_{3}(g) $$
Answer #1

molecules o gpeour eaetanmts nwo. of moles of parb So
Answer #2
Concepts and reason

Le Chatelier’s Principle: When a chemical reaction in equilibrium is changed (change in temperature, or pressure, or concentration of reactants, or products), the equilibrium will shift its direction to adjust the change to reestablish the equilibrium.

Fundamentals

Example for Le Chatelier’s principle:

A(g) + 2B(g)
C(g)+D(g)

In the above equilibrium reaction, the gaseous molecule in the reactant side is more than products. According to Le Chatelier’s principle, the equilibrium shifts its direction to adjust the change. Therefore, the position of equilibrium will move to the right (towards product). Therefore, the concentration of reactant decreases and concentration of product increases as the pressure increases.

Ng(8)+3H, (8) =2NH3(8)
Gaseous molecules in reactants = 4 (1+3)
Gaseous molecules in products = 2

N, (g)+3H2(g)=2NH, (g)
Down Down Up

Ans:

Concentration of each species after increasing the pressure is shown below.

N, (g)+3H2(g)=2NH, (g)
Down Down Up